OpenStudy (anonymous):

Can somebody explain why -64 squared is not a real number?

OpenStudy (anonymous):

squared or square root?

OpenStudy (anonymous):

-64^2 is 4096

OpenStudy (anonymous):

like, inside the squared box thing...which I don't know how to type..

OpenStudy (anonymous):

i^2 = -1 No rational number squared is (-64)

OpenStudy (anonymous):

\[\sqrt{-64}\]

OpenStudy (anonymous):

So it started out as -64 to the 1/2 power...

OpenStudy (anonymous):

I changed it to -64 inside the square root box thing..

OpenStudy (anonymous):

and my solutions manual says the answer is not a real number.

OpenStudy (anonymous):

if you take the square root of a negative number, it becomes imaginary

OpenStudy (anonymous):

something squared is something times itself. Nothing times itself will give u a negative number; u need a negative times a positive 4 that. So there's no such thing as the square root of -64, because nothing times itelf will give u a negative number

OpenStudy (anonymous):

You can take out the 8 and get: \[8\sqrt{-1}\] \[\sqrt{-1} = i\] So you get: \[8i\]

OpenStudy (anonymous):

\[\sqrt{-1}=i\]

OpenStudy (anonymous):

thanks